# How to Learn in a Noisy World? Self-Correcting the Real-World Data Noise on Machine Translation

### v1

@@ B.2 Misaligned Noise Samples
Table 6 shows the simualted misaligned samples of misaligned-laser and misaligned-comet. Overall, the simulated misaligned noise controlled by external models all share certain amounts of semantic meanings compared with the true reference.
@@
B.3 Human Evaluation of Misaligned Noise
The aim of this data annotation is to evaluate the adequacy of real-world and simulated misaligned noise. Given a source and misaligned target, the annotator should select the overlap meanings between them, as shown in Table 7. The simulated misaligned sentence pairs are constructed from the clean corpus WMT2017 DeEn and the real-world misaligned sentences are selected from web-mined Paracrawl datasets. The language pair we select here is German into English, and therefore we select the human annotator who are sufficient in both German and English languages.
@@ B.3 Human Evaluation of Misaligned Noise
Appendix C Training Details.
C.1 Training and Evaluation
We follow the setup of the Transformer-base and Transformer-big models (Bengio et al., 2015). For each model, the number of layers in the encoder and in the decoder is . For Transformer-base, we employ parallel attention layers or heads. The dimensionality of input and output is , and the inner layer of feed-forward networks has dimensionality . For Transformer-big, we employ parallel attention layers or heads. The dimensionality of input and output is , and the inner layer of feed-forward networks has dimensionality .
All models are trained with the Adam optimizer (Kingma and Ba, 2015) for up to 500K steps for high-resource tasks and 100K steps for low-resource tasks, with a learning rate of 5e-4 and an inverse square root scheduler. A dropout rate of 0.3 and label smoothing of 0.2 are used. Each model is trained on one NVIDIA A6000 GPU with a batch size of 25K tokens. We choose the best checkpoint according to the average validation loss of all language pairs. The data is tokenized with the SentencePiece tool (Kudo and Richardson, 2018) and we build a shared vocabulary of 32K tokens. For evaluation, we employ beam search decoding with a beam size of 5. BLEU scores are computed using detokenized case-sensitive SacreBLEU. 888nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1
C.2 Dataset Details
Appendix D Chrf++ and Comet Scores
@@
| dynamic | 35.8 | 34.4* | 33.6* | 35.8* | 31.5* | 28.3* | 35.8* | 35.0* | 33.4* | |
| COMET | ||||||||||
| enfr | enru | entr† | enes† | enbe† | enht† | ensi | enkm | Avg. | ||
| Misaligned Rate (%) | 10% | 19% | 44% | 22% | 10% | 35% | 62% | 18% | - | |
| Corpus Size (M) | 5M | 5M | 5M | 5M | 1.1M | 0.55M | 0.21M | 0.06M | - | |
| Baseline | 80.0 | 79.2 | 82.0 | 76.5 | 68.3 | 63.2 | 59.6 | 73.6 | 72.8 | |
@@
| dynamic | 81.6 | 80.6 | 83.0 | 77.9 | 68.5 | 63.6 | 63.3 | 75.0 | 74.2 | |
| ChrF++ | ||||||||||
| Baseline | 67.3 | 50.6 | 54.8 | 49.1 | 36.5 | 45.4 | 20.2 | 15.6 | 42.4 | |

### v2

@@ B.1 Algorithm
To save computational resources for calculating the LASER/COMET score for a source sentence with a chunk of target sentences, we first perform surface-level feature control (word overlap and length mismatch) to select a subset of misaligned target candidates. Word overlap is used as a filter to ensure that the misaligned targets share certain surface-level features with the true reference. Same for the length mismatch.
To avoid overusing the selected misaligned target, we remove the selected target from the chunk of target sentences . In our adequacy evaluation (shown in Appendix B.3 and Table 7), we also show that our misaligned sentences contain only partial meanings of the source sentences. This ensures a low likelihood that the selected misaligned target is a reasonable source sentence translation.
@@ B.2 Misaligned Noise Samples
Table 6 shows the simulated misaligned samples of Misaligned-LASER and Misaligned-COMET. Overall, the simulated misaligned noise controlled by external models all share certain amounts of semantic meanings compared with the true reference.
@@
B.3 Adequacy Evaluation
To evaluate the adequacy of the real world and our simulated misalignment noise, we design an annotation guide (see Table 7) to select the overlap meanings between a source sentence with the misaligned target. The simulated misaligned sentence pairs are constructed from the clean corpus WMT2017 DeEn, and the real-world misaligned sentences are selected from web-mined Paracrawl datasets. The annotations were conducted by the two PhD students, who are also the authors of this paper, as volunteers without compensation.
@@ B.3 Adequacy Evaluation
B.4 Token-level Annotation
We conducted a token-level annotation on 50 misaligned and clean sentences, resulting in 480 misaligned tokens and 1557 clean tokens. The annotators must label each token as “clean” or “noisy” given a source and a target sentence. The annotated misaligned and clean sentences are sampled from the sentences used in Section 3.3. The annotations were conducted by the two PhD students, who are also the authors of this paper, as volunteers without compensation.
Table 8 shows that the average el2n values for the misaligned tokens are higher than those for clean tokens, in both misaligned and clean samples, further confirming the effectiveness of leveraging the model’s self-knowledge to distinguish data noise. Moreover, we also find some clean tokens in clean target sentences do have higher el2n values (shown in Table 9). We find that clean tokens with higher el2n values tend to be difficult words for the model to learn, e.g., “communication” and “developments”.
| Misaligned | Clean-M | Clean-C |
| 1.13 | 0.32 | 0.37 |
| SRC | _ganz _entschieden _möchte _ich _mich _gegen _den _Ansatz _der _Kommission _wenden _, _wie _er _in _ihrer _Mitteilung _zum _Ausdruck _kommt _. |
| TGT | _I _should _also _like _to _firmly _contest _the _Commission _& apos ; s _approach _as _presented _in _its _communication _. |
| High el2n | _contest, _communication |
| SRC | _wir _werden _daher _diesen _Bericht _unterstützen _und _das _Thema _auch _weiterhin _mit _großer _Aufmerksamkeit _verfolgen _. |
| TGT | _we _therefore _support _this _report _and _will _continue _to _closely _monitor _developments _. |
| High el2n | _closely, _monitor, _developments |
| SRC | _folglich _muß _bis _zur _Revision _ein _ausreichen der _Zeitraum _ver gehen _. |
| TGT | _we _must _therefore _provide _for _a _review _after _a _sufficient _period _. |
| High el2n | _therefore |
Appendix C Self-Correction Method Design
C.1 Label Correction in Computer Vision
Our self-correction method is in line with the label correction method in Computer Vision (Wang et al., 2022; Lu and He, 2022). Both approaches are motivated by the idea of correcting data noise using a model’s self-knowledge. However, we are the first work to apply this approach specifically in the text de-noise field.
While other work (Kim et al., 2021) highlights another benefit of using the model’s predictions to refine the target, i.e. regularization. However, we do not discuss this aspect in our paper. This is because we share different motivations. Our work primarily aims to improve the robustness of training to address the low-quality training data issues instead of regularizing the model.
C.2 Hyper-Parameter Selection
In , decides the inflection point, and adjusts the exponentiation’s base and growth speed. Therefore, we fixed and conducted prior experiments to select . Table 10 provides the results of different under 30% misaligned noise rations for high-resource and low-resource tasks. We select and for our experiments.
| High-resource | Low-resource | |
| -4 | 31.9 | 19.7 |
| -5 | 32.0 | 20.0 |
| -6 | 32.3 | 20.3 |
| -7 | 31.8 | 20.2 |
| -8 | 31.4 | 20.0 |
C.3 The Impact of Sharpening Model Prediction.
Here, we aim to analyze the impact of sharpening model prediction distribution, i.e., different fixed values of , to correct the ground truth on translation performance. We train the self-correction models on DeEn task with 30% of different types of noise, with varying values of softmax temperature . From figure 5, we show that using sharpening model prediction distribution with a smaller achieves better translation performance for all noisy settings. However, the optimal varies when training with different types of noise and thus increases the difficulty of selecting a fixed for different scenarios. This motivates us to design a dynamic , which varies automatically in a low range of entropy state over training time. The overall performance in both Section 5.2 and Section 5.3 by using a dynamic also shows its general applicability for different noise scenarios.
Appendix D Training Details.
D.1 Training and Evaluation
We follow the setup of the Transformer-base and Transformer-big models (Bengio et al., 2015). For each model, the number of layers in the encoder and in the decoder is . We employ parallel attention layers and heads for the Transformer-base. The dimensionality of input and output is , and the inner layer of feed-forward networks has dimensionality . We employ parallel attention layers and heads for Transformer-big. The dimensionality of input and output is , and the inner layer of feed-forward networks has dimensionality .
All models are trained with the Adam optimizer (Kingma and Ba, 2015) for up to 500K steps for high-resource tasks and 100K steps for low-resource tasks, with a learning rate of 5e-4 and an inverse square root scheduler. A dropout rate of 0.3 and label smoothing of 0.2 are used. Each model is trained on one NVIDIA A6000 GPU with a batch size of 25K tokens. We choose the best checkpoint according to the average validation loss of all language pairs. The data is tokenized with the SentencePiece tool (Kudo and Richardson, 2018), and we build a shared vocabulary of 32K tokens. For evaluation, we employ beam search decoding with a beam size of 5. BLEU scores are computed using detokenized case-sensitive SacreBLEU999nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1.
D.2 Dataset Details
D.2.1 Simulated Noise Setting
Table 11 shows the training and evaluation dataset details for clean training corpus in simulated noisy experiments in Section 5.2.
@@ D.2.1 Simulated Noise Setting
D.2.2 Real-World Noise Setting
For Paracrawl, the language pairs are: enfr (French), ensi (Sinhala), ensw (Swahili), and enkm (Khmer). For CCAligned, the language pairs are entr (Turkish), enes (Spanish), and enbe (Belarusian). For the high-resource language pairs: enfr, entr, enes, we randomly sample 5M sentence pairs as the training corpus. For medium and low-resource language pairs, we use the original corpus size.
Appendix E Chrf++ and COMET Scores
Table 12, 13, and 14 shows the COMET (Unbabel/wmt22-comet-da) and Chrf++ scores for all experiments.
@@ Appendix E Chrf++ and COMET Scores
| dynamic | 35.8 | 34.4 | 33.6 | 35.8 | 31.5 | 28.3 | 35.8 | 35.0 | 33.4 | |
| COMET | |||||||||
| enfr | entr† | enes† | enbe† | ensi | ensw | enkm | Avg. | ||
| Misaligned Rate (%) | 10% | 44% | 22% | 10% | 62% | 11% | 18% | - | |
| Corpus Size (M) | 5M | 5M | 5M | 1.1M | 210K | 130K | 60K | - | |
| Baseline | 80.0* | 82.0* | 76.5* | 68.3* | 59.6* | 59.0* | 73.6* | 71.3 | |
@@
| dynamic | 81.6 | 83.0 | 77.9 | 68.9 | 63.6 | 61.4 | 75.0 | 73.0 | |
| ChrF++ | |||||||||
| Baseline | 67.3* | 54.8* | 49.1* | 36.5* | 20.2* | 37.9* | 15.6* | 40.2 | |